Table sugar (sucrose) molecules can be enzymatically broken down to produce fructose and dextrose (also called glucose). The reaction can be written

sucrose\overset{invertase.enzyme}{\rightarrow} dextrose + fructose

The following procedure was followed to determine the sucrose content of sweet potatoes. A 78.7-g sample of sweet potato was reduced to juice, and the juice was collected in a beaker. The jucing machine was washed 3x with 100 mL portions of sodium phosphate buffer with 2-3 min between rnsings. The juice and washings were poured into a 500-mL volumetric flask, and the containers were rinsed with several 10 mL portions of buffer, which were added to the flask. Buffer was then added "to the mark," which indicates the point where the volume is 500.0 mL. The sample was then refrigerated for 1 hour. After an hour, a 3mL aliquot of the sample solution was taken, and the invertase enzyme was added to it. The sample that was not enzymatically decomposed was assayed for free dextrose. The content was found to be 208 mg/dL. After 20 min, the sample with invertase was assayed and found to contain 458 mg/dL dextrose. In each case, a 25 uL sample is used.

a. How many mg of dextrose are due to the decomposition of sucrose in the sweet potato?

b. Given the molecular masses of dextrose (180.16) and sucrose (342.3), what is the percent sucrose (w/w) in the potatoes?
Answer #1

A) 1 dL= 100 mL
   Free dextrose =208 mg/dL = 208mg/ 100ml
Volume of inital sample was prepared = 500ml
     \therefore Free dextrose in the sample will be 1040mg/ 500ml
   i.e the amount of free dextrose in the intial potato sample was 1040mg in 78.7 g of sweet potato

After 20 min, the sample with invertase was assayed and found to contain 458 mg/dL dextrose.

the dextrose concentration was 458 mg/dL = 458mg/100ml
so similarly the dextrose after decompostion of sucrose will be 2290mg/500ml of solution
i.e the amount of total dextrose in the intial potato sample was 2290 mg in 78.7 g of sweet potato

\thereforemg of dextrose are due to the decomposition of sucrose in the sweet potato = (total dextrose-free dextrose)
   = (2290-1040) mg
   = 1250 mg of dextrose

B) From first calculation amount of hydrolysed dextrose is 1250 mg= 1.250 g
   number of moles of dextrose = weight of dextrose (in gm)/ molecular weight of dextrose (given = 180.16)
     number of moles of dextrose = 1.250 / 180.16 = 0.006938 moles
   As from the above reaction it has been seen that 1 molecule of sucrose results in 1 molecule of dextrose
   So, the number of mole of sucrose = number of moles of dextrose = 0.006938
     
Amount of sucrose = number of moles X molecular weight
   = 0.006938 X 342.3 = 2.37 g

  \thereforepercent sucrose (w/w) in the potatoes = (2.37/78.7) X 100 = 3.011 %